OpenStudy (anonymous):

Jose is paying for a $2 drink using bills in his wallet. He has four $1 bills, two $5 bills, and two $10 bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two $1 bills to pay for the drink? Show your work.

OpenStudy (anonymous):

4/8 * 3/7

OpenStudy (anonymous):

the $1 bills iare what's concerned here...4 out of 8 are $1...the first one he takes has a 1/2 shot at being a $1...the second has a 3/7 shot at being a We only care about the $1 bills...4 out of 8 are $1...the first one he takes has a 1/2 shot at being a $1...the second has a 3/7 shot at being a $1: so P = 1/2*3/7 p =3/14